OpenStudy (anonymous):

can someone help me with qs 4 please???

OpenStudy (anonymous):

\[\sqrt{4}?\]

OpenStudy (anonymous):

I don't know sorry

OpenStudy (anonymous):

you can find the actuallyy functions for them

OpenStudy (anonymous):

because you know f(x)

OpenStudy (anonymous):

F(x)=(x^2-5x) f(2x)=(4x^2-10x)=4x(x-2.5) now the intersection points will be 0 and 2.5 instead of 0,5

OpenStudy (anonymous):

for 3f(x) you have to multiply 3 with the ordinates of that points. Now you can draw it.

OpenStudy (anonymous):

the multiply by 3 would it be 3x^2-15x???

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

but that o and 5 again and in the mark scheme it says minimum points and there different to 0, 2.5 and 0,5

OpenStudy (anonymous):

im confused :S